# A new indeterminacy-based quantum theory

### v4

@@ 3 Why is EnDQT local?
However, we can go further in terms of presenting evidence for why CCMs are inappropriate and justify the rejection of the CMC that improves over the above one. A way of finding the limitations of the domain of applicability of the CCMs is by examining the more general models that putatively represent causal relations in the quantum domain, i.e., quantum causal models (QCMs). [^58] Costa & Shrapnel (2016), Allen et al. (2017), and Barrett et al. (2019). I will analyze how QCMs make some assumptions that CCMs don’t make, and that that these assumptions concern the quantum domain according to EnDQT. QCMs are, in principle, more general because they reduce to classical ones in a "classical limit." Like we found what is wrong with classical mechanics when we examine the more general theory, QT, which reduces to classical mechanics in some limit; we find what is wrong with the CCMs, when we adopt QCMs interpreted via EnDQT.
@@
As I will explain, QCMs will have the role of showing how EnDQT provides a local causal explanation of Bell-type correlations. Note that QCMs currently are only formulated for finite-dimensional Hilbert spaces. However, this isn’t as far as we can tell, in principle, a fundamental limitation. [^59] See Paunkovic & Vojinovic (2023) for an overview of possible challenges that need to be addressed. I will thus pose the following argument,
  • P-1) QCMs, which assume the quantum Markov condition (QMC) that is a generalization of the classical one, according to EnDQT, explain locally Bell-type/quantum correlations.
@@
  • P-2) Quantum causal models interpreted by EnDQT explicitly consider that systems that participate in those causal relations i) can assume indeterminate values represented via QT, ii) only assume determinate values when they interact with an SDC, and iii) where those relations are described via QT. i)-iii) are appropriate assumptions for EnDQT.
  • P-3) On the other hand, CCMs and the CMC, which arise in a specific limit from QCMs (Barrett et al., 2020), don’t make in general the same EnDQT-appropriate assumptions as QCMs.
  • C) Hence, according to EnDQT, CCMs, with their CMC, are inappropriate to provide an account of quantum causal relations contrary to QCMs. The latter provides an appropriate causal explanation of Bell-type/quantum correlations, which is local.
@@
Let’s turn to the justification of P-1). QCMs consider that each node in the causal DAG concerns a possible locus of interventions on the properties of a system. More concretely, each node is associated with a set of CP (completely positive) maps $\tau_{A_{1}}^{k_{A_{1}}\mid x_{A_{1}}}\otimes\ldots\otimes\tau_{A_{\mathrm{n}}}^{k_{A_{\mathrm{n}}}\mid x_{A_{\mathrm{n}}}},$ [^60] A quantum channel is a linear map $\varepsilon$ that is a completely positive trace preserving (CPTP) map. A map is a CPTP map if: a) it is trace-preserving, i.e., $\operatorname{Tr}(\rho)=\operatorname{Tr}(\varepsilon(\rho))$ for all density operators $\rho$, b) positive, i.e., $\varepsilon(\rho)\geq 0$ whenever the density operator $\rho\geq 0$, and c) completely positive. When only b) and c) are fulfilled, we have a completely positive (CP) map rather than a CPTP map. A CP-map can be associated with a positive operator-valued measure (POVM). See Nielsen & Chuang (2011). also called quantum instruments, instead of random variables as in the CCMs case. This set gives the "possibility space" that can be associated with the different ways the properties of a system with its quantum state can change under local interventions $x$, which correspond to the preparation of quantum systems, transformations, measurements on them, etc., each leading to different outcomes $k$.
@@
The QMC is defined through the causal DAG where the edges of the DAG are associated with quantum channels/completely positive trace-preserving (CPTP) maps. [^61] See the previous footnote. Examples of a quantum channel are unitary maps, evolution of the system with noise, etc. [^62] Each (quantum) node $A_{i}$ is associated with an input Hilbert space $\mathcal{H}_{A_{i}^{\text{input }}}$, which I will write as $A_{i}^{\text{input }}$, and an output Hilbert space $\mathcal{H}_{A_{i}^{\text{output }}}$, corresponding to the incoming and outgoing system, which I will write as $A_{i}^{\text{output }}$, and each edge is associated with an output Hilbert space of one node and the input Hilbert space of another node. When it is written $\rho_{B\mid DA}\rho_{C\mid AE}$, what is meant is that $\rho_{B\mid DA}\rho_{C\mid AE}=\rho_{B\mid DA}\otimes\rho_{C\mid AE}=$ $\left(\rho_{B\mid DA}\otimes I_{E^{\text{output }}}\otimes I_{C^{\text{input }}}\right)\left(\rho_{C\mid AE}\otimes I_{B^{\text{input }}}\otimes I_{D^{\text{output }}}\right)$, where $X^{\text{input }}$ and $X^{\text{output }}$ is the inputs and outputs of node X. Moreover, $\operatorname{Tr}_{A}\rho_{AB\mid C}=\rho_{B\mid C}$ and $\operatorname{Tr}_{B}\rho_{AB\mid C}=\rho_{A\mid C}$. Both CP and CPTP maps are written as positive semi-definite operators via the Choi-Jamiolkowski (CJ)-isomorphism. [^63] See, e.g., Barrett et al. (2019).
@@
The QMC representing a causal structure held fixed is written via the process operator $\sigma$, which is a CPTP map and factorizes analogously to the CMC. More precisely, a process operator $\sigma_{A_{1},\ldots,A_{n}}$ is compatible with a DAG G with nodes $A_{1},\ldots,A_{n}$, if and only if it obeys the quantum Markov condition (QMC, Barrett et al., 2019) where this condition says for all $i$, $l$ in the DAG G there are quantum channels such that $\left[\rho_{A_{i}\mid Pa\left(A_{i}\right)},\rho_{A_{l}\mid Pa\left(A_{l}\right)}\right]=0$, and
@@
We need to have $\left[\rho_{A_{i}\mid Pa\left(A_{i}\right)},\rho_{A_{l}\mid Pa\left(A_{l}\right)}\right]=0$ because the product of two positive operators is positive if and only if they commute. $\sigma_{A_{1},\ldots,A_{n}}$ factorize analogously to conditional probabilities in the CMC.
@@
An obstacle that one must face to provide a local causal explanation of Bell correlations via QCMs is to deal with their operationalism. Causal influences are typically understood by the possibility of "signaling" from one node to another. [^64] When all the relevant systems participating in causal relations are included (Barrett et al., 2019). The causal structure represented by QCMs represents the constraints on these signaling relations. So, node $X$ cannot signal to node $Y$ if and only if node $X$ doesn’t precede node $Y$ in the graph. Signaling between node $X$ and $Y$ can be understood as occurring when a variation in the choice of certain instruments/interventions performed at node $X$ can vary the probabilities of an outcome $\mathrm{k}$ concerning measurements performed at node $Y$.
One may worry that, like in other QTs such as Bohmian mechanics, [^65] See, e.g., Goldstein (2021). although there isn’t signaling, we still have non-local influences, and QCMs are hiding such influences. If we adopt EnDQT, which doesn’t consider that there are hidden non-local influences that cannot be used for signaling, we don’t need to have this worry. This is because, according to EnDQT, SDCs are necessarily involved in the influences that give rise to determinate values, and they represent local interactions between systems (see previous section). Furthermore, using the concept of signaling and an operationalist language is unnecessary, and we don’t need to adopt an account where signaling or causation is irreducible. We can rather consider that systems in a region influence the determinate value of certain systems in another region, where such influences are modally described/governed by QT, and QCMs allow us to represent and infer those influences.
Let’s see how adopting the point of view of EnDQT, QCMs provide a local causal (non-relationalist, non-retrocausal, and non-superdeterministic) [^66] See Wood & Spekkens (2015) for examples of non-local, superdeterministic, and retrocausal causal structures. common cause explanation of quantum correlations. Now, $A$, $B$, and $\Lambda$, represent spacetime regions, instead of classical variables. Consider below how, via the QMC and a version of the Born rule, we can represent the local common cause structure that explains Bell correlations (Figure 4),
@@
Note that eq.(11) is analogous to eq.(8). The systems prepared at the source act as common causes for Bell correlations, having indeterminate values until each system interacts with Alice and Bob’s measurement devices, giving rise to the correlated outcomes. $\rho_{\Lambda}$ via its subsystems represents the systems prepared at the source, which, for example, could be systems that have indeterminate values of spin-p (where p ranges over all possible directions of spin). We use $\rho_{\Lambda}$ to represent each system in the different regions separately by keeping track of the labels $A$ and $B$ and the channels $\rho_{B\mid\Lambda}$ and $\rho_{A\mid\Lambda}$. Each system evolves locally to region $A/B$, where Alice/Bob influences the outcomes that arise in $A$/$B$. This influence is represented via the quantum channel $\rho_{A\mid\Lambda}$ in the case of $A$, and $\rho_{B\mid\Lambda}$ in the case of $B$. $\rho_{A\mid\Lambda}$ and $\rho_{B\mid\Lambda}$ are identity channels that acting on the density operator $\rho_{\Lambda}$ representing the systems in region $\Lambda$, evolve them to regions $A$ and $B$, respectively. The influence that gives rise to the outcomes is also represented via the POVMs $\tau_{A}^{x\mid s\ SDC}$ in the case of Alice, where $s$ is her random measurement choice, and $x$ is her outcome/the determinate value of S, and analogously via $\tau_{B}^{y\mid t\ SDC}$ in the case of Bob. The superscript SDC means that these are interventions that give rise to a determinate value, connecting systems with an SDC, and correspond to other kinds of edges in the DAG in Fig. 4. Alice and Bob, due to their measurements, will lead the systems to become part of an SDC because they also belong to SDCs.
So, with the above account, EnDQT allows QCMs to be explicitly local and nonoperational. The local interactions at each wing are mediated by the SDCs, and these interactions, plus the prepared systems at the source, provide a local non-relational common cause explanation of quantum correlations.
@@
Note that by adopting EnDQT’s view of quantum states, it isn’t considered that the (local) measurement of Alice on the system affects the system of Bob and Bob, and vice-versa. We aren’t reifying quantum states. We can represent this situation via the following (what I will call) EnDQT-causal-DAG (Fig. 4), where the nodes in grey represent the systems that $don^{\prime}t$ belong to an SDC, and the arrows in grey represent their evolution and influences on the values of the systems these arrows point to. The nodes in black represent the systems that belong to an SDC (Alice/Bob). The arrows in black represent their interactions with other systems that give rise to the latter having determinate values, pointing to these systems. These arrows and interactions are mathematically represented by POVMs. EnDQT-causal-DAGs aim to highlight the fact that measurement-like interventions in QCMs involve systems that are locally connected with SDCs. [^67] Note that this DAG is different from the ones above involving the propagation of the DC.
@@
EnDQT approach to quantum causation is not restricted to Bell scenarios, but also has the benefit of being applicable to scenarios where it’s difficult to see how to apply QCMs coherently, such as in the popular extended Wigner’s friend-like scenarios. [^68] See, e.g., Bong et al. (2020); Brukner (2018), Frauchiger & Renner (2018), Ormrod et al. (2023), Myrvold (2002), and Schmid et al. (2023) for a review. Suppose we have two friends/agents in isolated space-like separated labs in each wing, [^69] I will focus on the scenario from Bong et al. (2020) and Brukner (2018). and one Wigner/agent next to each lab, where the friends share an entangled pair prepared at the source like in the Bell scenario. It’s also assumed that the lab is isolated in such a way that Wigner can perform arbitrary unitary operations on the contents of the lab. Here we have the case explained in section 2 of a process represented by pragmatic reversible decoherence models involving the target system and the friend or their measurement device. The friend is a macroscopic system that (supposedly) makes a measurement/decoheres her target system; however, Wigner can reverse this process.
@@
So, if we manage to achieve this isolation from the SDCs (see section 2), the Wigners can unitarily manipulate the friend plus her system, possibly reversing their state. We would then treat each friend and their target systems as being in an entangled superposition of states at each wing. Thus, the evolution of each system of the entangled pair to each wing and the "measurements" of each friend would be treated via a unitary channel that entangles each friend and their target system, where these channels would also be used to represent the causal structure of this situation. [^70] In a toy scenario, we could represent the quantum channels that entangle each friend in each wing by a CNOT gate in the CJ-form, and each friend as a being initially in the state $|0\rangle$. Then, as I have mentioned, the Wigners in each wing can unitarily manipulate or measure these entangled states. We could then calculate the probabilities for these measurement outcomes for the different measurement settings of the Wigners using a version of the Born rule like in eq.(10), providing a local common cause explanation for this situation. [^71] More on this in Pipa (forthcoming-b).
@@
Notice that, in case the above isolation from the SDCs is successful, contrary to what is assumed by the theorem underlying the scenario mentioned above, there is no joint probability distribution for the outcomes of the friends and Wigner. This is because the friends inside their labs don’t obtain any outcomes since they don’t interact with SDCs, which allows the Wigners to manipulate them unitarily. So, it rejects the so-called absoluteness of observed events assumption of this theorem, not because events aren’t absolute like relationalist views claim, but rather because the events concerning the outcomes of the friends don’t occur.
@@
One may wonder what the friend experiences when it is in a superposition? In other words, what’s going on with the friend if we isolate it (from interacting with SDCs). There are different possibilities: the friend could lack mental content. However, it might seem unsatisfactory or even hard to conceive what is for a friend to lack mental content. Instead, we can consider that friend-like systems when isolated from SDCs have indeterminate mental content, where this content depends on the indeterminate physical properties of their brain. When the lab is open, their indeterminate mental content becomes determinate. This possibility has the benefit of not being foreign to the philosophy of mind. Externalism about mental content is roughly the thesis that mental content depends on the external environment of the subject that has that mental content. [^72] Putnam (1975). The friend $having$ determinate mental content is dependent on an environment that renders that content determinate. [^73] See Appendix A for a response to an objection regarding the EnDQT’s account of Wigner’s friend scenarios and the friend’s experiences.

### v5

@@ 3 Why is EnDQT local?
There is at least one possible objection to this argument for rejecting the adequacy of CMC. This justification makes it unclear whether $some$ causal explanation of quantum correlations can be provided, and this deficiency could press us to reject other assumptions of Bell’s theorem instead of the CRCCP. In the end, one might think that CCMs are good inferential tools, and the problem is elsewhere. Taking into account this objection, we can develop a better argument for why the CMC and CCMs are inappropriate to give a causal account of Bell correlations according to EnDQT. I will call it, the argument for locality. This argument has the benefit of not necessarily appealing to the proof of Pearl and Verma, and we can use other considerations as well. The argument is based on the observation that a way of finding the limitations of the domain of applicability of the CCMs is by examining the more general models that putatively represent causal relations in the quantum domain, i.e., Quantum Causal Models (QCMs).909090[28, 3, 3]. Furthermore, we will analyze the limitations of CCMs by comparison to QCMs as interpreted by QT. More concretely, I will analyze how QCMs make some assumptions that CCMs don’t make and that these assumptions concern the quantum domain, according to EnDQT. QCMs are, in principle, more general because they reduce to classical ones in a “classical limit.” Like we found what is wrong with classical mechanics when we examined the more general theory, QT, which reduces to classical mechanics in some limit, we will find what is wrong with the CCMs when we adopt QCMs interpreted via EnDQT.
@@
As I will explain, QCMs will have the role of showing how EnDQT provides a local causal explanation of Bell-type/quantum correlations and how we can infer those relations. Note that QCMs currently are only formulated for finite-dimensional Hilbert spaces. However, this isn’t as far as we can tell, in principle, a fundamental limitation. [^91] See [77] for an overview of possible challenges that need to be addressed in order to extend QCMs to the infinite-dimensional case. I will thus pose the following argument that I will present via three core premises.
Let’s start with the first core premise of the argument:
@@
P1) A causal model accurately represents causal relations between quantum systems for EnDQT if and only if it can be understood as explicitly considering that systems that participate in those causal relations i) only assume determinate values with a certain probability given by the Born rule when they interact with an SDC, and ii) where those relations are described via QT, but without reifying the quantum states like EnDQT assumes. i) and ii) are what we will call EnDQT-appropriate assumptions.
Given CDC1)-CDC4), this premise is plausible. Systems have indeterminate values of any observable by default, only having determinate values with a certain probability given by the Born when they interact with members of an SDC, hence i). Furthermore, obviously, the causal model should describe those relations explicitly via QT as interpreted by EnDQT, so that it is clear that what is representing is what EnDQT aims to represent. This also requires adopting the EnDQT perspective on quantum states. Note that causal models, both classical and quantum, don’t wear their ontology on their sleeves. They require some interpretation. Their interpretation is theory-dependent to a certain degree, and so a theory-dependent argument has to be made as to whether they appropriately represent causal relations between quantum systems or not. If a causal model can be understood as providing i) and ii) explicitly, it should be seen as accurately representing causal relations according to EnDQT.
Let’s turn to turn to the second premise:
P2) If QCMs consider that i) systems only assume determinate values with a given Born probability when they interact with an SDC, and ii) where those relations are described via QT, but without reifying the quantum states like EnDQT assumes, then QCMs provide a local non-relational, non-superdeterministic, and non-retrocausal explanation of quantum correlations.
Note that this local explanation includes a common cause local explanation of Bell correlations. Showing that P2) is true will require an interpretation of QCMs in agreement with EnDQT. We will turn to that now.
@@
First, I will provide some minimal technical background. QCMs consider that each node in the causal DAG concerns a possible locus of interventions on the properties of a system. More concretely, each node is associated with a set of CP (completely positive) maps $\tau_{A_{1}}^{k_{A_{1}}\mid x_{A_{1}}}\otimes\ldots\otimes\tau_{A_{\mathrm{n}}}^{k_{A_{\mathrm{n}}}\mid x_{A_{\mathrm{n}}}},$ [^92] A quantum channel is a linear map $\varepsilon$ that is a completely positive trace preserving (CPTP) map. A map is a CPTP map if: a) it is trace-preserving, i.e., $\operatorname{Tr}(\rho)=\operatorname{Tr}(\varepsilon(\rho))$ for all density operators $\rho$, b) positive, i.e., $\varepsilon(\rho)\geq 0$ whenever the density operator $\rho\geq 0$, and c) completely positive. When only b) and c) are fulfilled, we have a completely positive (CP) map rather than a CPTP map. A CP-map can be associated with a positive operator-valued measure (POVM). See [73]. also called quantum instruments, instead of random variables as in the CCMs case. This set gives the “possibility space” that can be associated with the different ways the properties of a system with its quantum state can change under local interventions $x$, which correspond to the preparation of quantum systems, transformations, measurements on them, etc., each leading to different outcomes $k$.
@@
The QMC is defined through a causal DAG where the edges of the DAG are associated with quantum channels/completely positive trace-preserving (CPTP) maps. [^93] See the previous footnote. Examples of a quantum channel are unitary maps, evolution of the system with noise, etc. [^94] Each (quantum) node $A_{i}$ is associated with an input Hilbert space $\mathcal{H}_{A_{i}^{\text{input}}}$, written as $A_{i}^{\text{input}}$, and an output Hilbert space $\mathcal{H}_{A_{i}^{\text{output}}}$, written as $A_{i}^{\text{output}}$. Each edge is associated with the output Hilbert space of one node and the input Hilbert space of another node. When written $\rho_{B\mid DA}\rho_{C\mid AE}$, it means $\rho_{B\mid DA}\rho_{C\mid AE}=\rho_{B\mid DA}\otimes\rho_{C\mid AE}=(\rho_{B\mid DA}\otimes I_{E^{\text{output}}}\otimes I_{C^{\text{input}}})(\rho_{C\mid AE}\otimes I_{B^{\text{input}}}\otimes I_{D^{\text{output}}})$, where $X^{\text{input}}$ and $X^{\text{output}}$ are the inputs and outputs of node $X$. Moreover, $\operatorname{Tr}_{A}\rho_{AB\mid C}=\rho_{B\mid C}$ and $\operatorname{Tr}_{B}\rho_{AB\mid C}=\rho_{A\mid C}$. Both CP and CPTP maps are written as positive semi-definite operators via the Choi-Jamiolkowski (CJ)-isomorphism. [^95] See, e.g., [7].
@@
The QMC representing a causal structure held fixed is written via the process operator $\sigma$, which is a CPTP map and factorizes analogously to the CMC. More precisely, a process operator $\sigma_{A_{1},\ldots,A_{n}}$ is compatible with a DAG G with nodes $A_{1},\ldots,A_{n}$, if and only if it obeys the Quantum Markov Condition (QMC), [7]). This condition says for all $i$, $l$ in the DAG G there are quantum channels such that $\left[\rho_{A_{i}\mid Pa\left(A_{i}\right)},\rho_{A_{l}\mid Pa\left(A_{l}\right)}\right]=0$, and
@@
We need to have $\left[\rho_{A_{i}\mid Pa\left(A_{i}\right)},\rho_{A_{l}\mid Pa\left(A_{l}\right)}\right]=0$ because the product of two positive operators is positive if and only if they commute. $\sigma_{A_{1},\ldots,A_{n}}$ factorize, which leads them to be analogous to the conditional probabilities in the CMC.
@@
An obstacle that one must face to provide a local causal explanation of Bell correlations via QCMs is to deal with their operationalism. Causal influences are typically understood by the possibility of “signaling” from one node to another. [^96] When all the relevant systems participating in causal relations are included [7]. The causal structure represented by QCMs represents the constraints on these signaling relations. So, node $X$ cannot signal to node $Y$ if and only if node $X$ doesn’t precede node $Y$ in the DAG (e.g., see Figure 5, more on this below). Signaling between node $X$ and $Y$ can be understood as occurring when a variation in the choice of certain instruments/interventions performed at node $X$ can vary the probabilities of an outcome $\mathrm{k}$ concerning measurements performed at node $Y$.
However, one might worry that, as in other QTs such as Bohmian mechanics, [^97] See, e.g., [44]. although there is no signaling, there could still be non-local influences, and QCMs might be hiding such influences. If we adopt EnDQT, which doesn’t consider that there are hidden non-local influences that cannot be used for signaling, we don’t need to have this worry because systems involved in QCMs, according to EnDQT, only have determinate values when they interact with members of SDCs. So, SDCs are necessarily involved in these influences that give rise to determinate values, and they concern local interactions between systems (see Section 2).
Furthermore, since EnDQT does not require agents at the fundamental level, using the concept of signaling and an operationalist language is unnecessary for understanding what QCMs fundamentally are about. Moreover, we don’t need to adopt an account where signaling or causation is irreducible. We can rather consider that systems in a region influence the determinate value of certain systems in another region, where such influences are modally described/governed by QT, and QCMs allow us to represent and infer those influences.
Given this background, let’s now see how if QCMs consider that systems i) only assume determinate values with a certain Born probability when they interact with an SDC, and ii) where those relations are described via QT but without reifying the quantum states like EnDQT assumes, then QCMs provide a local non-relational, non-superdeterministic, and non-retrocausal explanation of quantum correlations. To do this, we will show how the truth of the antecedent leads to the truth of the consequent. Thus, we will have to interpret QCMs according to EnDQT and in agreement with i) and ii).
Now, $A$, $B$, and $\Lambda$, represent spacetime regions, instead of classical variables. Consider below how, via the QMC and a version of the Born rule, we can represent the local common cause structure that explains Bell correlations (Figure 5),
@@
Note that eq. (14) is analogous to eq.(11).
According to EnDQT, systems prepared at the source act as common causes for Bell correlations, having indeterminate values until each system interacts with Alice and Bob’s measurement devices, giving rise to the correlated outcomes/determinate values. $\rho_{\Lambda}$ via its subsystems represents the systems prepared at the source.
@@
To see this more concretely, let’s focus on the example of systems that have indeterminate values of spin-p (where p ranges over all possible directions of spin) in a Bell scenario with two parties sharing an entangled state. This example could, in principle, be extended to any finite-dimensional case. We use $\rho_{\Lambda}$ to represent each system in the different regions separately by keeping track of the labels $A$ and $B$ and the channels $\rho_{B\mid\Lambda}$ and $\rho_{A\mid\Lambda}$. Each system evolves locally to region $A/B$, where Alice/Bob influences the outcomes that arise in $A$/$B$. This influence is represented via the quantum channel $\rho_{A\mid\Lambda}$ in the case of $A$, and $\rho_{B\mid\Lambda}$ in the case of $B$. $\rho_{A\mid\Lambda}$ and $\rho_{B\mid\Lambda}$ are identity channels that acting on the density operator $\rho_{\Lambda}$ representing the systems in region $\Lambda$, evolve them to regions $A$ and $B$, respectively. [^98] Note that an identity channel leave the quantum state unchanged. The influence that gives rise to the outcomes/determinate values is also represented via the POVMs $\tau_{A}^{x\mid s\ SDC}$ in the case of Alice, where $s$ is her random measurement choice, and $x$ is her outcome/the determinate value of S, and analogously via $\tau_{B}^{y\mid t\ SDC}$ in the case of Bob. The superscript SDC means that these are interventions that give rise to a determinate value, connecting each one of the systems with an SDC, and correspond to other kinds of edges in the DAG in Figure 5. Alice and Bob, due to their measurements, will lead the systems to become part of an SDC because they also belong to SDCs. Importantly, the relations of influence represented via QCMs are represented via QT but without reifying the quantum states like EnDQT assumes. More concretely, by adopting EnDQT’s view of quantum states, it isn’t considered that the (local) measurement of Alice on the system affects the system of Bob and Bob, and vice-versa because we aren’t reifying quantum states, viewing them as representing by themselves certain causal relations. Other elements of EnDQT and QT that we are seeing here, with the $auxiliary$ support of quantum states, provide such representation. The Born rule in (14) is only applicable when we take into account the interventions that lead the target systems to belong to the SDCs in each lab.
@@
Let’s see how this works diagrammatically. We can represent the Bell situation via the following (what I will call) EnDQT-causal-DAG (Fig. 5), where the nodes in grey represent the systems that $don^{\prime}t$ belong to an SDC, and the arrows in grey represent their evolution and influences on the values of the systems these arrows point to. The nodes in black represent the systems that belong to an SDC (Alice/Bob). The arrows in black represent their interactions with other systems that give rise to the latter having determinate values, pointing to these systems. These arrows and interactions in grey are mathematically represented by POVMs. EnDQT-causal-DAGs aim to highlight the fact that measurement-like interventions in QCMs involve systems that are locally connected with SDCs. [^99] Note that this DAG is different from the ones above involving the propagation of the DC.
@@
Thus, we can see that the DAG in Figure 5 leads to a local common cause explanation of Bell correlations,100100100See [104] for examples of non-local, superdeterministic, and retrocausal causal structures. They differ from the one in Figure 5. which in principle could be extended to any quantum correlations (even the more tricky ones to explain, see below). The local interactions at each wing are mediated by the SDCs, and these interactions, plus the prepared systems at the source, provide a local non-relational, non-superdeterministic, non-retrocausal common cause explanation of quantum correlations, including Bell correlations. Therefore, EnDQT allows QCMs to be explicitly local and non-operational, and according to EnDQT, QCMs provide a local non-relational, non-superdeterministic, and non-retrocausal explanation of quantum correlations.
@@
To be clear, note that EnDQT’s approach to quantum causation is not restricted to Bell scenarios but also has the benefit of being applicable to scenarios where it is difficult to see how to apply QCMs coherently, such as in the popular extended Wigner’s friend-like scenarios.101101101See, e.g., [14, 17, 39, 76, 70, 90]. Suppose we have two friends/agents in isolated space-like separated labs in each wing,102102102I will focus on the scenario from [14, 16]. and one Wigner/agent next to each lab, where the friends share an entangled pair prepared at the source, as in the Bell scenario. It is also assumed that the lab is isolated in such a way that Wigner can perform arbitrary unitary operations on the contents of the lab. Here, we have the case explained in Section 2, which concerns a process represented by pragmatic reversible decoherence models involving the target system and the friend or their measurement device. The friend is a macroscopic system that (supposedly) makes a measurement/decoheres her target system; however, Wigner can reverse this process.
@@
So, if the isolation of the friend plus their system from the SDCs is achieved (see Section 2), the Wigners can unitarily manipulate the friend plus their system, possibly reversing their state. We would then treat each friend and their target systems at each wing as being in an entangled superposition of states. Thus, the evolution of each system of the entangled pair to each wing and the “measurements” of each friend would be treated via a unitary channel that entangles each friend and their target system, where these channels would also represent the causal structure of this situation.103103103In a toy scenario, we could represent the quantum channels that entangle each friend in each wing by a CNOT gate in the CJ-form, and each friend as a being initially in the state $|0\rangle$. The target systems would be in an entangled state in the computational basis, (i.e., a state belonging to a Hilbert space spanned by the basis $|00\rangle,|11\rangle,|10\rangle$ and $|01\rangle$). Then, as I have mentioned, the Wigners in each wing can unitarily manipulate or measure these entangled states. We could then calculate the probabilities for these measurement outcomes for the different measurement settings of the Wigners using a version of the Born rule like in eq. (13), providing a local common cause explanation for this situation.
@@
Notice that, in the (unlikely or perhaps even impossible) case that the above isolation from the SDCs is successful, contrary to what is assumed by the theorem underlying the scenario mentioned above, there is no joint probability distribution for the outcomes of the friends and Wigner. This is because the friends inside their labs don’t obtain any outcomes since they don’t interact with SDCs, which allows the Wigners to manipulate them unitarily. So, it rejects the so-called absoluteness of observed events assumption of this theorem, not because events aren’t absolute like relationalist views claim, but rather because the events concerning the outcomes of the friends don’t occur.104104104This assumption is typically explained in the following way: “[a]ny observed event is an absolute single event, and not relative to anything or anyone.” [14]. More precisely, there is a theoretical probability distribution $P$ of theory $\theta$ that can give rise to the empirical probability distribution $p$, where $p(a,b|x,y)=\sum_{c,d}P_{\theta}(a,b,c,d|x,y),\quad\forall a,b,x,y$ where $a$, $b$, $c$, and $d$ are the outcomes of Alice, Bob (which are the two Wigners, each in each wing), Charlie, and Debbie (which are the two friends, each within the isolated lab), respectively, and $x$ and $y$ are measurement choices of Alice and Bob. EnDQT is a theory that doesn’t consider that there is such theoretical probability distribution $P$ in the case the lab is isolated from the SDCs because Alice and Bob don’t obtain any outcomes. This theorem also assumes the locality assumption (explained in more detail below), which says that the outcomes of the system of one of the Wigners are statistically independent of the measurements of the other Wigner on their system and vice-versa. Furthermore, it assumes the no-superdeterminism assumption.
@@
Given its appeal to indeterminacy, one may wonder about what, according to EnDQT, the friend experiences when it is in a superposition. In other words, in the unlikely possibility that we manage to isolate such a macroscopic system from SDCs, what’s going on with their mental content (e.g., their thoughts, desires, etc.), according to EnDQT? More concretely, in the “local friendliness theorem” mentioned above,105105105[14]. Wigner sometimes opens the door and asks the friend which outcome they obtained; what’s happening with the mental content of the friend? There are different possibilities that I don’t have space to go in-depth here: one could consider that the friend lacks mental content. However, this position might seem unsatisfactory since it is hard to conceive what it is for a friend-like system to lack mental content. So, this possibility can be deemed as incoherent. Instead of adopting this position, I think that a more satisfactory possibility is to consider that friend-like systems, when isolated from SDCs, have indeterminate mental content, where this content depends on the indeterminate physical properties of their brain. When the lab is open, their indeterminate mental content becomes determinate.106106106See, e.g., [5] for a discussion of the relation between quantum theory and mental states. This possibility has the advantage of not being foreign to the philosophy of mind. Externalism about mental content roughly consists of the thesis that mental content depends on the external environment of the subject that has that mental content.107107107[84]. The friend $having$ determinate mental content is dependent on the SDCs that render that content determinate.
Due to its oddness, one may still object to this consequence of EnDQT, but I will now briefly argue that in actually experimentally attainable extended Wigner’s friend-like scenarios, many other respectable interpretations of QT will actually lead up to similar consequences. Extended Wigner’s friend scenarios might never be experimentally feasible because, arguably, we can never isolate macroscopic systems that are capable of measuring systems as the friend does in this scenario. However, recently [103] proposed a variation of this scenario where the Friend is an AI agent running on a quantum computer. Wigner is just a user of the quantum computer that is capable of manipulating the friend. Before Wigner erases the memory of the friend by unitarily manipulating this system, he can talk with the friend. This scenario makes extended Wigner’s friend-$like$ scenarios experimentally feasible in the future. They call it the local friendliness theorem. Granting that such a system is like us in some ways and, therefore, has mental content, one can similarly also consider that before the friend delivers their output to the exterior, their mental content is indeterminate. Thus, in this version, we would deny the “Friendliness” assumption used to derive this no-go theorem. Various influential interpretations of QT, such as the MWI and spontaneous collapse theory, would also deny this assumption. This is because for quantum computers to be functional, their internal states will need to be maintained in a coherent superposition in order to perform quantum computations. So, no collapse or branching should occur inside a quantum computer. Furthermore, in relationalist single-world theories, such as relational quantum mechanics,108108108[88, 32]. the friend would be considered to have indeterminate mental content.
Therefore, although one might still object to this feature of EnDQT, it is something that, in actually experimentally feasible Wigner’s friend scenarios, applies to other respectable interpretations of QT, such as collapse theories, MWI, and relational quantum mechanics. Thus, if we accept these other interpretations, we shouldn’t see this feature as a problem in our noisy worlds, where isolating macroscopic systems might never be feasible. In Appendix B, I will go into more detail about this argument.